OpenStudy (anonymous):

How to I Multiple this:

OpenStudy (anonymous):

\[\frac{ x ^{3}+1 }{ x ^{3}-x ^{2}+x }\times \frac{ 6x }{ -54x-54}\]

OpenStudy (anonymous):

(6 + x) * (-4 + x) x(6(-4 + x) + x(-4 + x)) = 24x x((-4 * 6 + x * 6) + x(-4 + x)) = 24x x((-24 + 6x) + x(-4 + x)) = 24x x(-24 + 6x + (-4 * x + x * x)) = 24x x(-24 + 6x + (-4x + x2)) = 24x

OpenStudy (anonymous):

.....What?

OpenStudy (anonymous):

yupp i am sure it is correct

OpenStudy (anonymous):

I don't even know what you did?

OpenStudy (anonymous):

Heeeelp?

OpenStudy (anonymous):

(6 + x) * (-4 + x) x(6(-4 + x) + x(-4 + x)) = 24x x((-4 * 6 + x * 6) + x(-4 + x)) = 24x x((-24 + 6x) + x(-4 + x)) = 24x x(-24 + 6x + (-4 * x + x * x)) = 24x x(-24 + 6x + (-4x + x2)) = 24x

OpenStudy (anonymous):

But what is that from? How did you do it all?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

So?